Jack & Joe leave their house at the same tme , drive towards each other.jack drives at60kmp and joe 30kmp.they pass each other in 1min.how far apart were Jack and joe when they started?
1
Expert's answer
2019-04-05T10:28:04-0400

The relative velocity is


"V=30+60=90 kmph"

The distance:


"S=Vt=(90 kmph)( \frac{1}{60} h)=1.5 km"